# XLB: A Differentiable Massively Parallel Lattice Boltzmann Library in Python

### v2

@@ Abstract
Open source software, Lattice Boltzmann Method, JAX, Machine learning, Differentiable programming, Scientific computing, Fluid simulation, Computational fluid dynamics, High performance computing
@@ 1 Introduction
In recent years, domain-specific libraries that are built on top of compiler technologies such as XLA, MLIR zhang2023compiler ; githubGitHubOpenxlaxla ; lattner2020mlir ; lattner2021mlir have gained substantial attention. These libraries mostly offer high-level programming interfaces while ensuring efficient execution by targeting specialized hardware like GPUs and TPUs. Libraries such as JAX jax2018github , PyTorch NEURIPS2019_9015 , Triton openaiIntroducingTriton , and TensorFlow tensorflow2015-whitepaper exemplify this approach. This trend is largely fueled by the increasing interest in machine learning and its various applications. Although predominantly employed in machine learning tasks, these libraries are also valuable for scientific computing and physics-based machine learning applications. They facilitate high-performance computation in Python, a language widely favored for its readability and ease of use, and provide tools for performing automatic differentiation, allowing for the application of machine learning within scientific domains. Numerous specialized libraries, such as JAX-Fluids bezgin2023jax , JAX-CFD kochkov2021machine (see also wang2022tensorflow ), BRAX brax2021github , JAX-MD schoenholz2020jax , JAX-FEM xue2023jax , Phiflow holl2020phiflow , TaichiDiff hu2019difftaichi and Taichi-LBM3D yang2022taichi , have been developed on top of these platforms.
@@
Recently, machine learning techniques have become increasingly recognized as valuable tool in scientific domains, ranging from molecular interactions and protein synthesis to astronomical and geophysical phenomena fluke2020surveying ; baron2019machine ; wang2018deepmd ; zhang2020dp ; jumper2021highly ; wang2022deep ; salehipour2019deep . Machine learning algorithms have been applied to many disciplines in physical sciences carleo2019machine ; brunton2021applying , to interpret large data sets, or uncover new scientific knowledge from raw data jia2021physics ; radovic2018machine ; vasudevan2021machine ; hochreiter2018machine . When combined with traditional scientific computing methods, machine learning can significantly enhance both the speed and accuracy of simulations and analyses edelen2020machine ; chen2022wavey . In fluid dynamics, the application of machine learning techniques has led to noble advancements in improving turbulence models, optimizing flow configurations, or predicting complex flow phenomena with high accuracy vinuesa2022enhancing ; kochkov2021machine ; fukami2020assessment ; salehipour2019deep . Differentiable fluid simulations when combined with deep learning approaches has demonstrated progress in a range of applications, from fast and accurate fluid flow prediction to learned turbulence models, shape optimization, and fluid control schenck2018spnets ; takahashi2021differentiable ; wandel2020learning ; belbute2020combining ; brahmachary2023unsteady ; holzschuh2023score ; um2020solver ; list2022learned ; chen2021numerical .
@@
The Lattice Boltzmann Method (LBM), originating from the kinetic gas theory, has emerged in recent decades as a widely accepted technique for tackling complex fluid dynamics problems. Its algorithmic foundation, built upon a simple yet effective ‘collide-and-stream’ mechanism on Cartesian grids, makes it highly parallelizable and exceptionally suitable for use with GPUs and TPUs. LBM has proven to be a promising methodology in simulating laminar and turbulent flows, subsonic and supersonic flows, flow through porous media, free-surface and multiphase flows, as well as complex flows involving fluid-solid interaction, combustion, and foaming phenomena huang2015multiphase ; renard2021improved ; liu2016multiphase ; grunau1993lattice ; mccracken2005multiple ; ataei2021lbfoam ; ataei2019numerical ; thurey2004free ; ginzburg2003lattice ; ataei2022hybrid ; schreiber2011free ; chiavazzo2009combustion . Recently, the integration of machine learning methods with LBM-based simulations has opened new frontiers in the field, providing enhanced capabilities such as the development of fast surrogate models and improved acoustic predictions ruttgers2020prediction ; chen2021compressed ; golsanami2022characterizing ; zhu2021numerical ; da2020ml ; hennigh2017lat ; bedrunka2021lettuce ; guo2016convolutional .
@@
In this paper, we introduce XLB, a differentiable massively parallel LBM library based on JAX, aimed to be used for large-scale fluid simulations, optimization, and physics-based machine learning. XLB has been developed with three key objectives: First, to prioritize accessibility, XLB employs Python and offers an interface that closely resembles Numpy, thereby ensuring ease-of-use and enabling quick adoption by a broad user base. Second, to emphasize extensibility, XLB adopts an object-oriented programming model, which allows users to effortlessly augment the library’s capabilities and tailor it to diverse research needs. Lastly, despite its user-friendly design, XLB does not sacrifice performance; it is engineered for high performance and scalability, making it suitable for both entry-level usage and advanced, resource-intensive applications.
@@
While open-source LBM libraries such as Palabos latt2021palabos and OpenLB krause2021openlb as well as high-performance programming models suitable for LBM calculations (see e.g. meneghin2022neon ) exist, the majority are written in low-level programming languages like C/C++. These libraries have high learning curves that make them less than ideal for rapid research and prototyping where the application itself may demand more focused investigation (rather than solver configuration and details). Hence, integration of these tools with available machine learning libraries such as PyTorch or JAX become complicated rendering the development of a unified computational model for physics-based machine learning very challenging. XLB has been designed to address these issues by providing an accessible and scalable solution for such applications. It successfully mitigates the performance shortfalls commonly associated with Python-based solutions, primarily through just-in-time (JIT) complication and distributed computing, positioning it as a compelling choice for research labs requiring a blend of high-performance and ease-of-use. Additionally, XLB’s adoption of the permissive Apache License, in contrast to the more restrictive GPL-based licenses of other LBM libraries, enhances its appeal for broader use in both academic and commercial applications.
@@
The most closely related software to XLB is Lettuce bedrunka2021lettuce , a PyTorch-based library that integrates LBM simulations with the PyTorch deep learning ecosystem. Nevertheless, XLB provides superior performance on single-GPU systems and offers the added advantage of being scalable on distributed, multi-GPU architectures. Furthermore, XLB’s intuitive and flexible programming model, derived from JAX’s Numpy-like interface, affords users a more straightforward path for extending the library’s functionality.
@@ 2 Lattice Boltzmann Method
The dynamics of an evolving flow field may be represented at mesoscopic scales using the LBM equations. These inherently time-dependent and discrete equations, govern the spatial and temporal behaviour of a set of velocity distribution functions or $f_{i}(x,t)$, through a collide-and-stream algorithm. In its most general form, the LBM equations may be formulated based on a general collision operator $\mathscr{C}$ as:
Collision: $f_{i}^{\ast}(x,t)=\mathscr{C}(f_{i}(x,t))$ (1)
Streaming: $f_{i}(x+\Delta x,t+\Delta t)=f_{i}^{\ast}(x,t)$ (2)
Notice that the subscript ‘$i$’ indexes the discrete lattice directions along which the above collide-and-stream operator applies. Each discrete cell situated at $x$, with spatial spacing $\Delta x=c_{i}\Delta t$, has a ‘lattice’ structure, formally denoted by a set of $q$ vectors $c_{i}=\{c_{1},\ldots,c_{q}\}$ that are visually illustrated in Figure 1 for 2D and 3D settings.
The collision operator in XLB may be of any form including (but not limited to) the single-relaxation model due to Bhatnagar-Gross-Krook (BGK), the multi-relaxation time (MRT) (see Kruger:2016:book ) method based on any form of the moment space, or the more advanced collision operators such as the cumulant collision Geier:2015:cumulant , recursive regularized coreixas2017recursive or the multi-relaxation entropic model (also known as the ‘KBC’ model) Karlin:2014:gibbs . Without loss of generality, we only present the classic BGK model here,
$\mathscr{C}_{\text{BGK}}:=f^{\ast}_{i}(x,t)=f_{i}(x,t)-\frac{\Delta t}{\tau}\left[f_{i}(x,t)-f_{i}^{eq}(x,t)\right]$ (3)
@@
in which $c_{s}$ is the constant speed of sound given by $c^{2}_{s}=(1/3)(\Delta x/\Delta t)^{2}$.
@@
where $w_{i}$ are weights associated with each lattice direction $c_{i}$. In addition, the macroscopic state variables namely the fluid density $\rho(x,t)$, velocity $u(x,t)$ and pressure $p(x,t)$ are formally derived from $f_{i}$ as,
density: $\rho(x,t)$ $=\sum_{i}f_{i}(x,t)$ (6)
velocity: $u(x,t)$ $=\frac{1}{\rho(x,t)}\sum_{i}c_{i}f_{i}(x,t)$ (7)
pressure: $p(x,t)$ $=c_{s}^{2}\rho(x,t)$ (8)
At the boundaries of the computational domain, where $x+c_{i}\Delta t$ would leave the assigned computational domain, it is necessary to impose boundary conditions such that the desired physical system can be simulated reasonably. Again there exists a multitude of options in the literature and all can be efficiently and easily implemented in XLB. For an updated list of available schemes in XLB the reader is referred to the online repository.
@@ 3 Implementation Details
3.1 Intro to JAX
JAX jax2018github , originally developed by Google, is a high-performance numerical computing library that offers significant advantages for scientific and engineering computations, particularly in the field of machine learning. It forms part of an extensive ecosystem of machine learning libraries such as Flax flax2020github , Equinox kidger2021equinox , and Optax deepmind2020jax , making it adaptable for various use cases.
@@ 3.1 Intro to JAX
The following sections will briefly explain the integration of JAX within XLB, highlighting how its features are leveraged to achieve high-performance computing in LBM.
While the XLB library is object-oriented, it is designed to be compatible with JAX, a library that emphasizes functional programming. JAX’s JIT compilation feature is generally incompatible with stateful functions, which are functions that exhibit side effects jax2018github . To reconcile this incompatibility with the object-oriented programming model used in XLB, functions that are jitted are decorated with partial that designates certain parameters (such as the self argument in classes) as static arguments, enabling JIT-compiled functions to operate within a object-oriented library, and still access static class attributes.
It is important to note that the XLB library is continually evolving, with frequent updates and enhancements. For the most up-to-date information, readers are advised to consult the XLB repository directly. As such, this section aims to provide a general overview of its structure and functionalities, considering that the specifics may evolve beyond the scope of this document at the time of the reader’s perusal.
@@ 3.4 Streaming Operation and Distributed Computing
The primary function, streaming_m, orchestrates the streaming processes and enables the inter-process communication in a multi-device context. This function isolates specific portions of populations from the left and right boundary slices of distribution arrays on each device. These isolated portions are then communicated to adjacent processes. streaming_p function handles streaming operations on a partitioned sub-domain within a single device. It leverages the vmap operation from the JAX library to carry out vectorized computations across all lattice directions. The streaming_i function executes individual streaming operations for each directional index $i$. This is achieved using the jnp.roll function from the JAX library, which shifts distribution function values in the direction specified by each discrete velocity vector $c_{i}$ (see Figure 1 for an illustration of these vectors for various lattice structures).
@@
To illustrate these streaming communications, Figure 3 provides a visual guide. First, the jnp.roll function shifts populations in the direction of discrete velocities $c_{i}$ (shown schematically for the D2Q9 lattice), handling the populations at the boundaries of each sub-domain in a toroidal manner. Subsequently, slices requiring communication to adjacent sub-domains are identified and permuted using send_right and send_left functions. The incoming data are assigned to the designated indices within the recipient sub-domain. This communication strategy effectively extends the toroidal shift behavior of jnp.roll across multiple devices, without requiring halo cells to be allocated for the multi-device communication.
@@ 3.5 Boundary Conditions
XLB includes methods for both setting up boundary attributes and applying boundary conditions (BCs). Specialized BCs inherit from this a base class and override methods and variables as needed for their specific implementation.
Certain BCs, like equilibrium and full-way bounce-back, merely require identification of the boundary cells for their application. In contrast, other BCs, such as Zou-He zou1997pressure , regularized Latt:2008:regularized or half-way bounce-back, demand additional supplementary data. This may entail knowing the types of the surrounding cells, the distance to the solid mesh, the direction of unknown populations requiring reconstruction by the BC scheme, the boundary’s normal vector, and conditions specifying whether the BC must be applied to the adjacent fluid cell. To facilitate this, an initialization step generates a boolean mask array similar to the size and cardinality of the population array $f_{i}$. In this array, for each direction $c_{i}$, the value is set to True if $-c_{i}$ points to solid neighbors (conceptually equivalent to directions where $f_{i}$ is streamed from solid neighbors) and conversely to False if $-c_{i}$ points to a fluid neighbor (conceptually equivalent to directions where $f_{i}$ is streamed from fluid neighbors).
@@
The BoundaryCondition class uses the mask array to modify the allocated boundary cells appropriately (if required by the boundary type) and to generate any supplementary information required for implementing the BC scheme such as the normal direction or the momentum flux.

### v3

@@ Abstract
Open source software , Lattice Boltzmann Method , JAX , Machine learning , Differentiable programming , Scientific computing , Computational fluid dynamics , High performance computing
journal: Computer Physics Communications: https://doi.org/10.1016/j.cpc.2024.109187
@@ 1 Introduction
In recent years, domain-specific libraries that are built on top of compiler technologies such as XLA and MLIR [1, 2, 3, 4] have gained substantial attention. These libraries mostly offer high-level programming interfaces while ensuring efficient execution by targeting specialized hardware like GPUs and TPUs. Libraries such as JAX [5], PyTorch [6], Triton [7], and TensorFlow [8] exemplify this approach. This trend is largely fueled by the increasing interest in machine learning and its various applications. Although predominantly employed in machine learning tasks, these libraries are also valuable for scientific computing and physics-based machine learning applications. They facilitate high-performance computation in Python, a language widely favored for its readability and ease of use, and provide tools for performing automatic differentiation, allowing for the application of machine learning within scientific domains. Numerous specialized libraries, such as JAX-Fluids [9], JAX-CFD [10] (see also [11]), BRAX [12], JAX-MD [13], JAX-FEM [14], Phiflow [15], TaichiDiff [16] and Taichi-LBM3D [17], have been developed on top of these platforms.
@@
Recently, machine learning techniques have become increasingly recognized as valuable tool in scientific domains, ranging from molecular interactions and protein synthesis to astronomical and geophysical phenomena [18, 19, 20, 21, 22, 23, 24]. Machine learning algorithms have been applied to many disciplines in physical sciences [25, 26], to interpret large data sets, or uncover new scientific knowledge from raw data [27, 28, 29, 30]. When combined with traditional scientific computing methods, machine learning can significantly enhance both the speed and accuracy of simulations and analyses [31, 32]. In fluid dynamics, the application of machine learning techniques has led to novel advancements in improving turbulence models, unsteady flow prediction, optimizing flow configurations, or predicting complex flow phenomena with high accuracy [33, 34, 10, 35, 24]. Differentiable fluid simulations when combined with deep learning approaches have demonstrated progress in a range of applications, from fast and accurate fluid flow prediction to learned turbulence models, shape optimization, and fluid control [36, 37, 38, 39, 40, 41, 42, 43, 44].
@@
The Lattice Boltzmann Method (LBM), originating from the kinetic gas theory, has emerged in recent decades as a widely accepted technique for tackling complex fluid dynamics problems. Its algorithmic foundation, built upon a simple yet effective ‘collide-and-stream’ mechanism on Cartesian grids, makes it highly parallelizable and exceptionally suitable for use with GPUs and TPUs. LBM has proven to be a promising methodology in simulating laminar and turbulent flows, subsonic and supersonic flows, flow through porous media, free-surface and multiphase flows, as well as complex flows involving fluid-solid interaction, combustion, and foaming phenomena [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. Recently, the integration of machine learning methods with LBM-based simulations has opened new frontiers in the field, providing enhanced capabilities such as the development of fast surrogate models and improved acoustic predictions [57, 58, 59, 60, 61, 62, 63, 64].
@@
In this paper, we introduce XLB, a differentiable massively parallel LBM library based on JAX, aimed to be used for large-scale fluid simulations, optimization, and physics-based machine learning. XLB has been developed with three key objectives. First, to prioritize accessibility, XLB employs Python and offers an interface that closely resembles Numpy, thereby ensuring ease-of-use and enabling quick adoption by a broad user base. Second, to emphasize extensibility, XLB adopts an object-oriented programming model, which allows users to effortlessly augment the library’s capabilities and tailor it to diverse research needs. Lastly, despite its user-friendly design, XLB does not sacrifice performance; it is engineered for high performance and scalability, making it suitable for both entry-level usage and advanced, resource-intensive applications.
@@
While open-source LBM libraries such as Palabos [65] and OpenLB [66] as well as high-performance programming models suitable for LBM calculations (see e.g. [67]) exist, the majority are written in low-level programming languages like C/C++. These libraries have high learning curves that make them less than ideal for rapid research and prototyping where the application itself may demand more focused investigation (rather than solver configuration and details). Hence, integration of these tools with available machine learning libraries such as PyTorch or JAX becomes complicated rendering the development of a unified computational model for physics-based machine learning very challenging. XLB has been designed to address these issues by providing an accessible and scalable solution for such applications. It successfully mitigates the performance shortfalls commonly associated with Python-based solutions, primarily through just-in-time (JIT) compilation and distributed computing, positioning it as a compelling choice for research labs requiring a blend of high-performance and ease-of-use. Additionally, XLB’s adoption of the permissive Apache License, in contrast to the more restrictive GPL-based licenses of other LBM libraries, enhances its appeal for broader use in both academic and commercial applications.
@@
The most closely related software to XLB is Lettuce [63], a PyTorch-based library that integrates LBM simulations with the PyTorch deep learning ecosystem. Nevertheless, XLB provides superior performance and offers the added advantage of being scalable on distributed multi-GPU architectures. Furthermore, XLB’s intuitive and flexible programming model, derived from JAX’s Numpy-like interface, affords users a more straightforward path for extending the library’s functionality.
@@ 2 Lattice Boltzmann method
The dynamics of an evolving flow field may be represented at mesoscopic scales using the LBM equations. These inherently time-dependent and discrete equations, govern the spatial and temporal behaviour of a set of velocity distribution functions or $f_{i}(\bm{x},t)$, through a collide-and-stream algorithm. In its most general form, the LBM equations may be formulated based on a general collision operator $\mathscr{C}$ as:
Collision: $f_{i}^{\ast}(\bm{x},t)=\mathscr{C}(f_{i}(\bm{x},t))$ (1)
Streaming: $f_{i}(\bm{x}+\Delta\bm{x},t+\Delta t)=f_{i}^{\ast}(\bm{x},t)$ (2)
Notice that the subscript ‘$i$’ indexes the discrete lattice directions along which the above collide-and-stream operator applies. Each discrete cell situated at $\bm{x}$, with spatial spacing $\Delta\bm{x}=\bm{c}_{i}\Delta t$, has a ‘lattice’ structure, formally denoted by a set of $q$ vectors $\bm{c}_{i}=\{\bm{c}_{1},\ldots,\bm{c}_{q}\}$ that are visually illustrated in Figure 1 for 2D and 3D settings.
The collision operator in XLB may be of any form including (but not limited to) the single-relaxation model due to Bhatnagar-Gross-Krook (BGK), the multi-relaxation time (MRT) method (see [68]) based on any form of the moment space, or the more advanced collision operators such as the cumulant collision [69], recursive regularized [70] or the multi-relaxation entropic model (also known as the ‘KBC’ model) [71]. Without loss of generality, we only present the classic BGK model here,
$\mathscr{C}_{\text{BGK}}:=f^{\ast}_{i}(\bm{x},t)=f_{i}(\bm{x},t)-\frac{\Delta t}{\tau}\left[f_{i}(\bm{x},t)-f_{i}^{eq}(\bm{x},t)\right]$ (3)
@@
in which $c_{s}$ is the constant speed of sound given by $c^{2}_{s}=(1/3)(\Delta\bm{x}/\Delta t)^{2}$.
@@
where $w_{i}$ are weights associated with each lattice direction $\bm{c}_{i}$. In addition, the macroscopic state variables namely the fluid density $\rho(\bm{x},t)$, velocity $\bm{u}(\bm{x},t)$ and pressure $p(\bm{x},t)$ are formally derived from $f_{i}$ as,
density: $\rho(\bm{x},t)$ $=\sum_{i}f_{i}(\bm{x},t)$ (6)
velocity: $\bm{u}(\bm{x},t)$ $=\frac{1}{\rho(\bm{x},t)}\sum_{i}\bm{c}_{i}f_{i}(\bm{x},t)$ (7)
pressure: $p(\bm{x},t)$ $=c_{s}^{2}\rho(\bm{x},t)$ (8)
At the boundaries of the computational domain, where $\bm{x}+\bm{c}_{i}\Delta t$ would leave the assigned computational domain, it is necessary to impose boundary conditions such that the desired physical system can be simulated reasonably (see Section 3.5 for further details). Again there exists a multitude of options in the literature and all can be efficiently and easily implemented in XLB. For an updated list of available schemes in XLB the reader is referred to the online repository.
@@ 3 Implementation details
The following sections will briefly explain the integration of JAX within XLB, highlighting how its features are leveraged to achieve high-performance computing in LBM. It is important to note that the XLB library is continually evolving, with frequent updates and enhancements. For the most up-to-date information, readers are advised to consult the XLB repository directly. As such, this section aims to provide a general overview of its structure and functionalities, considering that the specifics may evolve beyond the scope of this document at the time of the reader’s perusal.
3.1 A brief introduction to JAX
JAX [5], originally developed by Google, is a high-performance numerical computing library that offers significant advantages for scientific and engineering computations, particularly in the field of machine learning. It forms part of an extensive ecosystem of machine learning libraries such as Flax [72], Equinox [73], and Optax [74], making it adaptable for various use cases.
@@ 3.1 A brief introduction to JAX
XLB incorporates object-oriented design for certain functionalities. To ensure compatibility with JAX, a functional programming library, XLB leverages the partial decorator. This decorator freezes specific parameters (like the self argument in classes) as static during JIT compilation process. This allows JIT-compiled functions to work within XLB’s object-oriented structure while avoiding the side effects that arise from JAX’s preference for stateless functions [5].
@@ 3.3 Simulation domain and array representation
Our approach shards XLB arrays across multiple GPUs, enabling global operations on the entire dataset without requiring separate computations. While efficient and easy to use, some operations may still benefit from explicit communication between GPUs. JAX provides protocols for this targeted communications, allowing us to fine-tune data distribution and processing during complex distributed operations. Section 3.4 of the manuscript details their use for the streaming operation.
@@ 3.4 Streaming operation and distributed computing
The primary function, streaming_m, orchestrates the streaming processes and enables the inter-process communication in a multi-device context. This function isolates specific portions of populations from the left and right boundary slices of distribution arrays on each device. These isolated portions are then communicated to adjacent processes. streaming_p function handles streaming operations on a partitioned sub-domain within a single device. It leverages the vmap operation from the JAX library to carry out vectorized computations across all lattice directions. The streaming_i function executes individual streaming operations for each directional index $i$. This is achieved using the jnp.roll function from the JAX library, which shifts distribution function values in the direction specified by each discrete velocity vector $\bm{c}_{i}$ (see Figure 1 for an illustration of these vectors for various lattice structures).
@@
To illustrate these streaming communications, Figure 3 provides a visual guide. First, the jnp.roll function shifts populations in the direction of discrete velocities $\bm{c}_{i}$ (shown schematically for the D2Q9 lattice), handling the populations at the boundaries of each sub-domain in a toroidal manner. Subsequently, slices requiring communication to adjacent sub-domains are identified and permuted using send_right and send_left functions. The incoming data are assigned to the designated indices within the recipient sub-domain. This communication strategy effectively extends the toroidal shift behavior of jnp.roll across multiple devices, without requiring halo cells to be allocated for the multi-device communication.
@@ 3.5 Boundary conditions
XLB includes methods for both setting up boundary attributes and applying boundary conditions (BCs). Specialized BCs inherit from a base class and override methods and variables as needed for their specific implementation.
Certain BCs, like equilibrium and full-way bounce-back, merely require identification of the boundary cells for their application. In contrast, other BCs, such as Zou-He [75], regularized [76] or half-way bounce-back, demand additional supplementary data. This may entail knowing the types of the surrounding cells, the distance to the solid mesh, the direction of unknown populations requiring reconstruction by the BC scheme, the boundary’s normal vector, and conditions specifying whether the BC must be applied to the adjacent fluid cell. To facilitate this, an initialization step generates a boolean mask array similar to the size and cardinality of the population array $f_{i}$. In this array, for each direction $\bm{c}_{i}$, the value is set to True if $-\bm{c}_{i}$ points to solid neighbors (conceptually equivalent to directions where $f_{i}$ is streamed from solid neighbors) and conversely to False if $-\bm{c}_{i}$ points to a fluid neighbor (conceptually equivalent to directions where $f_{i}$ is streamed from fluid neighbors).
@@
The boundary conditions rely on the mask array to modify the allocated boundary cells appropriately (if required by the boundary type) and to generate any supplementary information required for implementing the BC scheme such as the normal direction or the momentum flux.